OpenStudy (anonymous):

2(theta)cos(THETA)+THETA=0

OpenStudy (dumbcow):

factor out a theta \[\rightarrow \theta (2\cos(\theta) +1) = 0\]

OpenStudy (anonymous):

how would you solve it? even after factoring out? littile confused?

OpenStudy (dumbcow):

solve by setting each factor equal to 0 \[\theta = 0\] \[2 \cos \theta +1 = 0\]

OpenStudy (anonymous):

2cosθ+1=0 =1+sin2θ

OpenStudy (anonymous):

^?

OpenStudy (anonymous):

2cosθ+1=0 2cosθ=-1 cosθ=-1/2